A gardener is planting flowers in a circular bed. If he plants 1256 flowers which are p.lanted 4 to a square metre,find the diameter of this bed
Answer by jim_thompson5910(35256)

"he plants 1256 flowers which are planted 4 to a square metre", so 1256/4 = 314 square meters are used up.


Area of a circle
A = pi*r^2

Plug in the area A = 314 and solve for r

A = pi*r^2
314 = pi*r^2 ... replace A with 314
314 = 3.14*r^2 ... replace pi with 3.14
314/3.14 = (3.14*r^2)/3.14 ... Divide both sides by 3.14.
100 = r^2
r^2 = 100
sqrt(r^2) = sqrt(100) ... take the square root of both sides
r = 10

So the radius is 10, making the diameter 2*10 = 20 meters